RhondaSommer (rhondasommer):

Evaluate [3 - (3 - 7(4 - 7) + 7)] + 2

OpenStudy (alexandervonhumboldt2):

first of all we need to solve what is in parentheses: so what is 4-7=?

OpenStudy (witty_roger):

BODMAS rule Bracket Of Division Multiplication Addition Subtraction

OpenStudy (alexandervonhumboldt2):

-6 is wrong

RhondaSommer (rhondasommer):

so sorry, my little brother...

RhondaSommer (rhondasommer):

-3

OpenStudy (phi):

Each step is not too hard, but you need to do them in the correct order start with [3 - (3 - 7(4 - 7) + 7)] + 2 notice you have (4-7) in parens. do that first, 4-7 is -3 (as you know) replace the (4-7) with -3, and you now have \[ [ 3 - (3 -7\cdot -3 +7)]+2 \] you still have some parens. look at (3- 7*-3 +7) you have add/subtract and multiply which do you do first ?

RhondaSommer (rhondasommer):

multiply

RhondaSommer (rhondasommer):

pemdas

RhondaSommer (rhondasommer):

@phi

OpenStudy (phi):

yes, so in (3- 7*-3 +7) what do you multiply and what do you get ?

RhondaSommer (rhondasommer):

3+21+7 = 31 ?

OpenStudy (phi):

yes, but you should be sure of how you get that. one way: 7*-3 is -21 so you now do (3 - (-21) + 7) = 3+21 +7 2nd way write it as (3 + -7*-3 + 7) notice instead of 3 -7 we can write it as 3+ -7 (it means the same thing) now do -7*-3 to get 21 and you have (3+21+7)

RhondaSommer (rhondasommer):

yay i did it right :) :)

OpenStudy (phi):

minus signs can be a pain, so try to learn the tricks. meanwhile [3 - 31] + 2 now do the parens (or brackets in this case, which they used for "readability")

RhondaSommer (rhondasommer):

-26?

OpenStudy (phi):

3-31 is -(31-3) (the rule is do big minus small, and use the sign of the bigger number) so -28 now add 2 -28+2 I would just make -21 two bigger to get -26 or you can write it 2+-28 or 2-28 and use the rule -(28-2) = -26

RhondaSommer (rhondasommer):

i kept adding wrong so i was getting confused. then it didnt help that i entered it wrong in the calculator :P

RhondaSommer (rhondasommer):

thanks for the help :)

OpenStudy (phi):

yes, arithmetic can require "attention to detail". If you have a tendency to lose your place or make silly mistakes then you have to slow down and think through each step. Like anything, (for example , learning to read or write), you will get it with practice. (Of course you have to work hard.)

RhondaSommer (rhondasommer):

im great at reading and wrting and usually these kind of equations. i dont know what happened :P would it be okay to tag you if i have other math questions?

OpenStudy (phi):

yes

OpenStudy (witty_roger):

sorry..its 30